# The relativistic rotated harmonic oscillator

### v1

@@ 1 Introduction
The motivation for the present study is not merely our mathematical curiosity, but also the new concept of unconventional representation of quantum observables by non-self-adjoint operators [32, 26]. In particular, it follows from our results that $H_{\theta}$ is not quasi-self-adjoint (i.e. similar, via a bounded and boundedly invertible transformation, to a self-adjoint operator). More generally, despite the growing interest in non-self-adjoint Dirac operators, very few explicitly solvable models are available (see [5] from the list above for an exception). In this paper, we offer the community a relativistic analogue of the highly influential Davies’s oscillator (1.4), which involves both the illusory simplicity due to the availability of explicit solutions as well as all the non-self-adjoint peculiarities.
Let us now present our main results. First of all, we realise $H_{\theta}$ as a closed operator with compact resolvent (see Proposition 3.2). Contrary to the sectorial Schrödinger case (1.4), however, the numerical range of $H_{\theta}$ coincides with the whole complex plane whenever $\theta\not=0$. The crucial observation is the supersymmetric relationship
@@ 2 Preliminaries
In this section, we are particularly concerned with the definition of the operator $H_{\theta}$ from (1.7).
@@ 2.1 The Dirac matrices
are the Pauli matrices. The Dirac (as well as Pauli) matrices are Hermitian and the anticommutation relations $\{\alpha_{\mu},\alpha_{\nu}\}=2\delta_{\mu\nu}$ hold with every $\mu,\nu\in\{0,1,2,3\}$. These anticommutation properties lead to (1.8), where the diagonal matrix explicitly reads
@@ 2.2 The non-relativistic harmonic oscillator
$\mathsf{D}({S_{\theta}}):=H^{2}(\mathbb{R})\cap L^{2}(\mathbb{R},x^{2}\,\mathrm{d}x)\,,$ (2.4)
@@
$\mathcal{D}:=H^{1}(\mathbb{R})\cap L^{2}(\mathbb{R},x\,\mathrm{d}x)\,.$ (2.5)
@@
(The natural numbers $\mathbb{N}$ contain zero in our convention.) The eigenvalues are (algebraically) simple.
@@
In particular, the wild behaviour (1.10) follows. Moreover, it is known that this inclusion is optimal in the sense that $|\vartheta|=|\theta|$ is indeed the transition angle for the pseudospectral properties of $S_{\theta}$. Indeed,
@@ 3.1 Definition
with a positive constant $K$, where we have used the boundedness of $\Omega$. Hence, since $\Psi\upharpoonright\Omega\subset L^{2}(\Omega)^{4}$, the above inequality shows that $\Psi\upharpoonright\Omega\in H^{1}(\Omega)^{4}$ and therefore $\Psi\in H^{1}_{\mathrm{loc}}(\mathbb{R})^{4}$.
@@
We continue now by an argument analogous to the proof of [33, Thm. 4.3]. Let $f\in C^{\infty}_{0}(\mathbb{R})$ be such that $f(x)=1$ for $|x|\leq 1$ and let $f_{n}(x)=f(x/n)$ with $n\geq 1$. From the elliptic regularity we have that $\Psi\in L^{2}(\mathbb{R})^{4}$ and $(A+i)\Psi=0$ implies that $\Psi\in C^{\infty}(\mathbb{R})$. It follows that:
$(A+i)(f_{n}\Psi)(x)=-if_{n}^{\prime}(x)\alpha_{1}\Psi(x)=-\frac{i}{n}f^{\prime}(x/n)\alpha_{1}\Psi(x)$
@@
These conditions are equivalent to $|\theta|<\frac{\pi}{2}$ and $k^{2}>1$, respectively. By [22, Thm. IV.1.16], it follows that $A+ik+B$ with $\mathsf{D}(A+B):=\mathsf{D}(A)=\mathcal{D}^{4}$ is closed and bijective. Moreover, the inverse $(A+ik+B)^{-1}$ is compact. Since adding $ik$ as well as $m\alpha^{3}$ represent just bounded perturbations, we conclude with the following result (where $m$ is possibly non-zero).
@@ 4.2 The Dirac operator
$=\sqrt{2(n+1)}\,\phi_{n+1}(\hat{x})=\sqrt{2(n+1)}\,\varphi_{n+1}(x)$
@@ 4.3 The asymptotic behaviour of the eigenprojectors
For the non-relativistic harmonic operator $S_{\theta}$, it is easy to see that $\|P_{n}\|=\|\varphi_{n}\|\,\|\tilde{\varphi}_{n}\|$. Moreover, $\|\varphi_{n}\|=\|\tilde{\varphi}_{n}\|$. Then (2.8) follows by the asymoptotics of Hermite functions.
@@ 4.4 The non-relativistic limit
${\hat{S}}_{\theta}=\frac{1}{2m}\left(-e^{-i\theta}\partial^{2}_{x}+m^{2}\omega^{2}e^{i\theta}x^{2}\right)\,{,}\qquad\mathsf{D}({\hat{S}}_{\theta}){:=\mathsf{D}(S_{\theta})=}H^{2}(\mathbb{R})\cap L^{2}(\mathbb{R},x^{2}\,\mathrm{d}x)\,.$
@@ Acknowledgment
D.K. was supported by the EXPRO grant No. 20-17749X of the Czech Science Foundation. A.B. and J.M.P.P. acknowledge support provided by the “Agencia Estatal de Investigación (AEI)” Research Project PID2020-117477GB-I00, by the QUITEMAD Project P2018/TCS-4342 funded by the Madrid Government (Comunidad de Madrid-Spain) and by the Madrid Government (Comunidad de Madrid-Spain) under the Multiannual Agreement with UC3M in the line of “Research Funds for Beatriz Galindo Fellowships” (C&QIG-BG-CM-UC3M), and in the context of the V PRICIT (Regional Programme of Research and Technological Innovation). A.B. acknowledges financial support from the Spanish Ministry of Universities through the UC3M Margarita Salas 2021-2023 program (“Convocatoria de la Universidad Carlos III de Madrid de Ayudas para la recualificación del sistema universitario español para 2021-2023”).

### v2

@@ 1 Introduction
The motivation for the present study is not merely our mathematical curiosity, but also the concept of unconventional representation of quantum observables by non-self-adjoint operators. This observation goes back to a 1992 article by physicists [32] concerned with bounded operators similar to self-adjoint ones. However, there are important mathematical intricacies for unbounded operators, which can be conveniently handled by the notion of pseudospectra, see [26] for a survey.
It follows from the present results that $H_{\theta}$ is actually not quasi-self-adjoint (i.e. similar, via a bounded and boundedly invertible transformation, to a self-adjoint operator). Our argument is again based on a pseudospectral analysis, namely, on studying the growth of the norm of the resolvent for large complex energies.
More generally, despite the growing interest in non-self-adjoint Dirac operators, very few explicitly solvable models are available (see [5] from the list above for an exception). In this paper, we offer the community a relativistic analogue of the highly influential Davies’s oscillator (1.4), which involves both the illusory simplicity due to the availability of explicit solutions as well as all the non-self-adjoint peculiarities.
Let us now present our main results. First of all, we realise $H_{\theta}$ as a closed operator with compact resolvent (see Theorem 3.2). Contrary to the sectorial Schrödinger case (1.4), however, the numerical range of $H_{\theta}$ coincides with the whole complex plane whenever $\theta\not=0$. The crucial observation is the supersymmetric relationship
@@ 2 Preliminaries
In this section we are going to introduce notation and some results that we shall need for the definition of the operator $H_{\theta}$ from (1.7).
@@ 2.1 The Dirac matrices
are the Pauli matrices. The Dirac (as well as Pauli) matrices are Hermitian and the anticommutation relations $\{\alpha_{\mu},\alpha_{\nu}\}:=\alpha_{\mu}\alpha_{\nu}+\alpha_{\nu}\alpha_{\mu}=2\delta_{\mu\nu}I_{\mathbb{C}^{4}}$ hold with every $\mu,\nu\in\{0,1,2,3\}$. These anticommutation properties lead to (1.8), where the matrix $i\alpha_{1}\alpha_{2}$ reads
@@ 2.2 The non-relativistic harmonic oscillator
$\mathsf{D}({S_{\theta}}):=H^{2}(\mathbb{R})\cap L^{2}(\mathbb{R},x^{4}\,\mathrm{d}x)\,,$ (2.4)
@@
$\mathcal{D}:=H^{1}(\mathbb{R})\cap L^{2}(\mathbb{R},x^{2}\,\mathrm{d}x)\,.$ (2.5)
@@
(The set of natural numbers $\mathbb{N}$ contains zero in our convention.) The eigenvalues are (algebraically) simple.
@@
In particular, the wild behaviour (1.10) follows. Moreover, it is known that this inclusion is optimal in the sense that $|\vartheta|=|\theta|$ is the transition angle for the pseudospectral properties of $S_{\theta}$. Indeed,
@@ 3.1 Definition
with a positive constant $K$, where we have used the boundedness of $\Omega$. Hence, since $\Psi\upharpoonright\Omega\in L^{2}(\Omega)^{4}$, the above inequality shows that $\Psi\upharpoonright\Omega\in H^{1}(\Omega)^{4}$ and therefore $\Psi\in H^{1}_{\mathrm{loc}}(\mathbb{R})^{4}$.
@@
We continue now by an argument analogous to the proof of [33, Thm. 4.3]. Let $f\in C^{\infty}_{0}(\mathbb{R})$ be such that $f(x)=1$ for $|x|\leq 1$ and let $f_{n}(x)=f(x/n)$ with $n\geq 1$. From the elliptic regularity we have that $\Psi\in L^{2}(\mathbb{R})^{4}$ and $(A+i)\Psi=0$ imply that $\Psi\in C^{\infty}(\mathbb{R})^{4}$. It follows that:
$(A+i)(f_{n}\Psi)(x)=-if_{n}^{\prime}(x)\alpha_{1}\Psi(x)=-\frac{i}{n}f^{\prime}(x/n)\alpha_{1}\Psi(x).$
@@
These conditions are equivalent to $|\theta|<\frac{\pi}{2}$ and $k^{2}>1$, respectively. By [22, Thm. IV.1.16], it follows that $A+ik+B$ with $\mathsf{D}(A+B):=\mathsf{D}(A)=\mathcal{D}^{4}$ is closed and bijective with compact inverse. Since adding $ik$ as well as $m\alpha^{3}$ represent just bounded perturbations, we conclude with the following result (where $m$ is possibly non-zero).
@@ 4.2 The Dirac operator
$=\sqrt{2(n+1)}\,\phi_{n+1}(\hat{x})=\sqrt{2(n+1)}\,\varphi_{n+1}(x).$
@@ 4.3 The asymptotic behaviour of the eigenprojectors
For the non-relativistic harmonic operator $S_{\theta}$, it is easy to see that $\|P_{n}\|=\|\varphi_{n}\|\,\|\tilde{\varphi}_{n}\|$, where $P_{n}$ are the projections associated to the non-relativistic rotated harmonic oscillator, cf. Eq. (2.8). Moreover, $\|\varphi_{n}\|=\|\tilde{\varphi}_{n}\|$. Then (2.8) follows by the asymoptotics of Hermite functions.
@@ 4.4 The non-relativistic limit
${\hat{S}}_{\theta}=\frac{1}{2m}\left(-e^{-i\theta}\partial^{2}_{x}+m^{2}\omega^{2}e^{i\theta}x^{2}\right)\,{,}\qquad\mathsf{D}({\hat{S}}_{\theta}){:=\mathsf{D}(S_{\theta})=}H^{2}(\mathbb{R})\cap L^{2}(\mathbb{R},x^{4}\,\mathrm{d}x)\,.$
@@ Acknowledgment
D.K. was supported by the EXPRO grant No. 20-17749X of the Czech Science Foundation. A.B. and J.M.P.P. acknowledge support provided by the “Agencia Estatal de Investigación (AEI)” Research Project PID2020-117477GB-I00, by the QUITEMAD Project TEC-2024/COM-84-QUITEMAD-CM funded by the Madrid Government (Comunidad de Madrid-Spain) and by the Madrid Government (Comunidad de Madrid-Spain) under the Multiannual Agreement with UC3M in the line of “Research Funds for Beatriz Galindo Fellowships” (C&QIG-BG-CM-UC3M), and in the context of the V PRICIT (Regional Programme of Research and Technological Innovation). A.B. acknowledges financial support from the Spanish Ministry of Universities through the UC3M Margarita Salas 2021-2023 program (“Convocatoria de la Universidad Carlos III de Madrid de Ayudas para la recualificación del sistema universitario español para 2021-2023”).